Find the domain and the range of the function. 4 f(x)=1-7x+ 5, for x < 7. for x 27 What is the domain of the given function? (Type your answer in interval notation.) What is the range of the graph of a function? (Type your answer in interval notation.) Enter your answer in each of the answer boxes
Answer #1

Clearly, f(x) is defined for every value of x<7 and x>=7.

So, f(x) is defined for all real numbers.

So, Domain of the function is R where R is a set of real numbers.

In interval-notation Domain is (-infinty,infinity) i.e oc.. oc .

We can get range by putting x= -\infty ,

f(x)= -\infty

for x= 7 , f(x) = 1

So, range is from -\infty to 1

In set interval notation, Range is (-\infty,1)